Name: 1. The time required for a citizen to complete the 2000 U.S. Census long form is normally distributed with a mean of
Answer #1

1.

Filling time, X ~ Normal(μ, σ)

μ = 40
σ = 10

P(X < 60) = P{(X - μ)/σ < (60 - 40)/10}

= P(Z < 2)

= 0.9773 [use table]

So, 0.97.73% of the citizens will require less than 60 minutes.

2.

The slowest 10% is at the right-most side of the distribution.

So, F(Z) = 1 - 0.1 = 0.90

or, Z = 1.28 [use table]

So, X = μ + Z.σ = 40 + 1.28*10 = 52.8

So, the slowest 10% will need 52.8 minutes or more for filling the form.

3.

(b) Statistics

4.

(d) All of the above

5.

σ = 0.10 cm
n = 12

So, Std. Error (σ) = σ / √n = 0.10/√12 = 0.029

6.

μ = 1.0
σ = 0.029

P(0.99 < x̄ < 1.01)

= P(x̄ < 1.01) - P(x̄ < 0.99)

= P(z < (1.01 - 1)/0.029) - P(z < (0.99 - 1)/0.029)

= P(z < 0.345) - P(z < -0.345)

= 0.635 - 0.365 [use table]

= 0.270